# TeLL-Drive: Enhancing Autonomous Driving with Teacher LLM-Guided Deep Reinforcement Learning

### v1

@@ Abstract
Although Deep Reinforcement Learning (DRL) and Large Language Models (LLMs) each show promise in addressing decision-making challenges in autonomous driving, DRL often suffers from high sample complexity, while LLMs have difficulty ensuring real-time decision making. To address these limitations, we propose TeLL-Drive, a hybrid framework that integrates an Teacher LLM to guide an attention-based Student DRL policy. By incorporating risk metrics, historical scenario retrieval, and domain heuristics into context-rich prompts, the LLM produces high-level driving strategies through chain-of-thought reasoning. A self-attention mechanism then fuses these strategies with the DRL agent’s exploration, accelerating policy convergence and boosting robustness across diverse driving conditions. Our experimental results, evaluated across multiple traffic scenarios, show that TeLL-Drive outperforms existing baseline methods, including other LLM-based approaches, in terms of success rates, average returns, and real-time feasibility. Ablation studies underscore the importance of each model component, especially the synergy between the attention mechanism and LLM-driven guidance. These findings suggest that TeLL-Drive significantly enhances both the adaptability and safety of autonomous driving systems, while offering a more efficient and scalable approach for policy learning. Full validation results are available on Our Website.
@@ I Introduction
Autonomous driving technology has made significant advancements over the past decade, emerging as a transformative force poised to revolutionize the transportation sector[1]. By promising enhanced safety, reduced traffic congestion, and increased mobility accessibility, autonomous vehicles (AVs) are set to redefine the landscape of modern transportation. Central to the operational efficacy of AVs is their ability to perform real-time, complex decision-making that rivals or surpasses human driving capabilities. Achieving such sophisticated decision-making necessitates the integration of advanced artificial intelligence methodologies capable of perceiving, interpreting, and responding to dynamic and often unpredictable driving environments[2].
Deep Reinforcement Learning (DRL) has become a key framework for decision-making in autonomous systems [3, 4]. In autonomous driving, DRL is used to develop policies for vehicle behaviors, such as intersection navigation [5]. However, traditional DRL faces challenges, including high data requirements, slow convergence, and limited environmental understanding [6], hindering its scalability and efficiency in dynamic driving scenarios that demand robust adaptability.
@@
Concurrently, LLMs, exemplified by architectures such as GPT-4, have demonstrated exceptional proficiency in understanding and generating human-like text. Leveraging vast repositories of knowledge and advanced contextual reasoning capabilities, LLMs have been applied to decision-making processes in autonomous driving[7, 8, 9]. However, their practical deployment as standalone decision-making agents is constrained by inherent challenges[10]. Specifically, LLMs struggle to ensure real-time responsiveness and exhibit a degree of randomness in their decision outputs, which are critical limitations in time-sensitive and safety-critical applications inherent to autonomous driving systems.
@@
To address these challenges, we propose TeLL-Drive, a framework that synergistically combines the strengths of DRL and LLMs to enhance decision-making in autonomous vehicles. TeLL-Drive mitigates the inherent limitations of DRL and LLMs by leveraging the contextual understanding and reasoning capabilities of LLMs to improve the sampling efficiency and quality of the DRL process. Specifically, we develop a risk-aware LLM agent equipped with memory, reflection, and reasoning abilities, enabling efficient and safe decision-making in complex traffic environments. Meanwhile, a DRL agent based on the Actor-Critic architecture is designed, which employs hybrid strategies to enhance sampling efficiency and quality while preserving robust exploration capabilities.
@@
As shown in Fig. 1, LLMs function as ”teachers”, providing high-quality guidance and contextual insights that inform and streamline the learning process of DRL agents. Subsequently, DRL serves as the ”student”, acting as the final decision maker to ensure real-time responsiveness and mitigate the randomness associated with LLM-driven decisions. The main contributions of this article are listed as follows:
@@
    TeLL-Drive is proposed, a novel framework for autonomous driving decision-making that combines a “teacher” LLM with a “student” DRL agent, which integrates the LLM’s high-level reasoning and knowledge with DRL’s adaptability and computational efficiency.
@@
    Through experiments in multiple scenarios, TeLL-Drive outperforms standard DRL algorithms in exploration efficiency and achieves favorable overall results compared to alternative methods.
@@ II-A DRL for Autonomous Driving Decisions
DRL has shown strong potential in autonomous driving, spanning tasks from lane-keeping to complex multi-agent interactions [11, 12]. However, two main challenges remain. Firstly, DRL often requires extensive environment interactions, which is costly and time-consuming. Secondly, lack of transparency impedes reliable decision-making in rare or out-of-distribution scenarios. To mitigate these issues, Reinforcement Learning from Human Feedback (RLHF) [13] has been proposed to incorporate expert knowledge, accelerating convergence and improving policy robustness. Nonetheless, RLHF remains costly due to the extensive human annotations required [14], and its feedback may not cover all possible driving conditions.
@@ II-B LLMs in Decision-Making
LLMs have exhibited significant promise in various tasks, including high-level reasoning for autonomous driving [15]. For instance, LanguageMPC [16] leverages an LLM’s common-sense reasoning to guide Model Predictive Control parameters, while Fu et al. [17] and Wen et al. [18] demonstrate LLM-based reasoning, interpretation, and memory capabilities for driving decisions. Despite these advances, direct deployment of LLMs faces obstacles. While high computational overhead makes LLMs difficult to achieve rapid, consistent responses required for safe driving, stochastic text generation may lead to unpredictable actions, which is risky in safety-critical environments.
II-C Hybrid RL-LLM Approaches
There is growing interest in combining RL and LLMs, though most efforts focus on using RL to refine LLMs rather than leveraging LLMs to enhance RL—particularly for autonomous driving. Initial studies have shown the potential of LLM-based reasoning to improve exploration efficiency and learning effectiveness of RL[19]. For example, Zhang et al. [20] introduced a semi-parametric RL framework with LLM-driven long-term memory, Trirat et al. [21] employed LLMs for automated machine learning, and Ma et al. [22] automated RL reward function design. However, fully exploiting LLMs for environmental understanding and decision-making in autonomous driving remains an open challenge, as existing approaches lack a systematic integration strategy that fully harnesses LLMs’ knowledge and reasoning capabilities.
@@ III-2 Action space
[…]
@@ IV-A Framework overview
TeLL-Drive leverages the knowledge of LLMs to guide the exploration and learning of DRL agents in various traffic scenarios. By introducing policy mixing, TeLL-Drive enhances sample efficiency and optimizes learning outcomes. As illustrated in Fig. 2, the framework comprises two main components: the LLM Teacher and the DRL Student. Based on multi-module collaboration, the Teacher Agent generates robust decision in complex traffic scenarios. The Teacher Agent generates robust decisions through its three key modules: Decision Engine, which provides real-time guidance; Memory Repository, which stores past experiences for context; and Reflective Evaluator, which refines the guidance based on previous performance. While the Student Agent refines the Teacher’s actions through a multi-head attention-based policy-mixing mechanism, integrating its own exploration experiences to effectively acquire knowledge from the LLM and enhance learning efficiency and quality.
@@ IV-B1 Decision Engine
Building on these real-time and historical insights, the engine constructs a comprehensive prompt $\mathcal{P}_{t}$ that integrates TTCP-derived risks, scenario-specific experiences, and traffic knowledge. This prompt includes road geometry, vehicle positions, and conflict zones, along with domain heuristics to provide the LLM with a rich contextual foundation for action proposals. To enhance reliability and minimize hallucinations, we then employ a chain-of-thought approach in which the model iteratively evaluates collision severity, short- and long-term maneuver consequences, and broader traffic implications. This structured reasoning process reduces logical inconsistencies, resulting in safer and more interpretable autonomous driving policies.
@@ IV-B2 Memory Repository
$\mathcal{P}_{t}\leftarrow\text{Retrieval}\Bigl{(}\mathbf{s}_{t},\mathcal{M}\Bigr{)}.$ (5)
@@ IV-B3 Reflective Evaluator
Any episode with $\max_{t}\Omega(s_{t},a_{t})\geq\delta$ is flagged for further reflection.
@@ IV-C1 Actor-Critic Algorithm with Policy Constraint
An actor-critic framework is adopted, where both the state-value function $V^{\pi}$ and the action-value function $Q^{\pi}$ are recursively estimated. For a policy $\pi(\mathbf{a}\mid\mathbf{s})$, the state-value function and the corresponding action-value function at state $\mathbf{s}_{t}$ is:
@@ IV-C2 Policy Distillation and Fusion
$\hat{\pi}^{T}(\mathbf{a}\mid\mathbf{s}_{t})$ $=\text{softmax}\bigl{(}\mathcal{W}_{p}\,\mathbf{h}_{t}^{T}+\mathbf{b}_{p}\bigr{)},$ (14)
$\hat{Q}^{T}(\mathbf{s}_{t},\mathbf{a})$ $=\mathcal{W}_{q}\,\mathbf{h}_{t}^{T}+\mathbf{b}_{q},$ (15)
@@
$\mathbf{h}_{t}=\alpha_{t}\,V+\mathbf{h}_{t}^{S}=\alpha_{t}\,\mathcal{W}_{V}\,\mathbf{h}_{t}^{T}+\mathbf{h}_{t}^{S}.$ (18)
@@
$\tilde{\pi}(\mathbf{a}\mid\mathbf{s}_{t})$ $=softmax\bigl{(}\mathbf{W}_{\pi}\,\mathbf{h}_{t}+\mathbf{b}_{\pi}\bigr{)},$ (19)
$\tilde{Q}(\mathbf{s}_{t},\mathbf{a})$ $=\mathbf{W}_{Q}\,\mathbf{h}_{t}+\mathbf{b}_{Q},$ (20)
@@
$\mathcal{L}_{\mathrm{distill}}=-\,\mathbb{E}_{(s,\mathbf{a}_{T})\in\mathcal{D}_{T}}\bigl{[}\log\,\hat{\pi}^{T}(\mathbf{a}_{T}\mid s)\bigr{]}.$ (21)
@@
V Experiments and Performance Evaluation
@@ V-B Implementation Details
Both our model and the baseline methods utilize a policy network composed of a multilayer perceptron (MLP) with two hidden layers of size $128\times 128$. We employ two self-attention heads, each also of dimension 128, to fuse the Student and Teacher representations. We set the discount factor $\gamma=0.99$ and adopt a learning rate of $5\times 10^{-4}$. Each model is trained for at least $10^{5}$ time steps, with an evaluation performed every 500 time steps. We use GPT-4o-mini as our LLM backbone, which shows reliable logical reasoning and real-time decision-making in driving tasks; it serves as the Teacher Agent for only the first 10% of training steps, after which constraints are gradually relaxed to encourage independent exploration. All experiments run on a computing platform equipped with Intel(R) Core(TM) i7-14700K CPU, an NVIDIA GeForce RTX 4080 SUPER GPU and 32 GB of RAM.

### v3

@@ Abstract
Although Deep Reinforcement Learning (DRL) and Large Language Models (LLMs) each show promise in addressing decision-making challenges in autonomous driving, DRL often suffers from high sample complexity, while LLMs have difficulty ensuring real-time decision making. To address these limitations, we propose TeLL-Drive, a hybrid framework that integrates a Teacher LLM to guide an attention-based Student DRL policy. By incorporating risk metrics, historical scenario retrieval, and domain heuristics into context-rich prompts, the LLM produces high-level driving strategies through chain-of-thought reasoning. A self-attention mechanism then fuses these strategies with the DRL agent’s exploration, accelerating policy convergence and boosting robustness across diverse driving conditions. The experimental results, evaluated across multiple traffic scenarios, show that TeLL-Drive outperforms existing baseline methods, including other LLM-based approaches, in terms of success rates, average returns, and real-time feasibility. Ablation studies underscore the importance of each model component, especially the synergy between the attention mechanism and LLM-driven guidance. Finally, we build a virtual-real fusion experimental platform to verify the real-time performance, robustness, and reliability of the algorithm running on real vehicles through vehicle-in-loop experiments. Full validation results are available on Our Website.
Index Terms:
Autonomous Vehicle; Large Language Model; Deep Reinforcement Learning
@@ I Introduction
Autonomous driving technology has made significant advancements over the past decade, emerging as a transformative force poised to revolutionize the transportation sector[1, 2]. By promising enhanced safety, reduced traffic congestion, and increased mobility accessibility, autonomous vehicles (AVs) are set to redefine the landscape of modern transportation. Central to the operational efficacy of AVs is their ability to perform real-time, complex decision-making that rivals or surpasses human driving capabilities. Achieving such sophisticated decision-making necessitates the integration of advanced artificial intelligence methodologies capable of perceiving, interpreting, and responding to dynamic and often unpredictable driving environments[3].
Deep Reinforcement Learning (DRL) has emerged as a key framework for autonomous decision-making[4, 5], with its ability to develop policies for complex tasks such as navigation through intersections [6, 7] and ramp merging [8, 9]. DRL’s strength lies in its capacity to learn from experience and optimize driving policies based on trial and error. However, despite its potential, traditional DRL methods face several challenges, including high data demands, slow convergence rates, and limited generalization across diverse and dynamic driving environments [10]. These limitations hinder the scalability and efficiency of DRL in real-world autonomous driving tasks, where adaptability, safety, and real-time performance are paramount.
@@
In parallel, Large Language Models (LLMs), exemplified by architectures such as GPT-4o[11], have demonstrated exceptional proficiency in natural language understanding and contextual reasoning. Leveraging vast repositories of knowledge and advanced contextual reasoning capabilities, LLMs can provide valuable insights for decision-making processes in autonomous driving systems[12, 13, 14, 15]. However, the deployment of LLMs as standalone decision-making agents faces significant barriers[13]. Specifically, LLMs struggle to ensure real-time responsiveness and exhibit a degree of randomness in their decision outputs, which are critical limitations in time-sensitive and safety-critical applications inherent to autonomous driving systems.
@@
To address the limitations, we propose TeLL-Drive, a novel framework that synergistically combines the strengths of both DRL and LLMs to enhance decision-making in autonomous vehicles. By leveraging the contextual understanding and reasoning capabilities of LLMs, TeLL-Drive enhances the sampling efficiency and quality of DRL, while mitigating the data inefficiency and slow convergence typically associated with DRL. Specifically, we introduce a risk-aware LLM agent, equipped with memory, reflective, and reasoning capabilities, that provides context-sensitive guidance to the DRL agent. This enables safer, more efficient decision-making in complex and dynamic traffic scenarios. Meanwhile, the DRL agent, built on the Actor-Critic architecture, ensures robust exploration and real-time responsiveness by employing hybrid strategies, addressing the inherent randomness in LLM-driven decisions.
@@
As shown in Fig. 1, the LLM function as a teacher, providing expert-level guidance and contextual insights that inform and streamline the learning process of DRL agents. Subsequently, DRL serves as the “student”, acting as the final decision maker to ensure real-time responsiveness and mitigate the randomness associated with LLM-driven decisions. The main contributions of this article are listed as follows:
@@
    We introduce TeLL-Drive, a decision-making framework for autonomous driving that combines a teacher LLM with a student DRL agent, which integrates the LLM’s high-level reasoning and knowledge with DRL’s adaptability and computational efficiency.
@@
    Through real-vehicle experiments in multiple scenarios, TeLL-Drive outperforms standard DRL algorithms in exploration efficiency and achieves favorable overall results compared to alternative methods.
@@ II-A DRL for Autonomous Driving Decisions
DRL has emerged as a promising approach for autonomous driving, spanning tasks from basic lane-keeping to complex multi-agent interactions [16, 17]. DRL algorithms, both policy-gradient-based and value-based, have demonstrated substantial performance improvements in simulated driving environments [18, 19]. These methods, by learning from interactions with the environment, can develop highly effective policies for tasks such as intersection navigation [6], obstacle avoidance, and adaptive cruise control. However, two main challenges persist in applying DRL to autonomous driving.
First, DRL’s reliance on extensive environment interactions often leads to high data requirements, which can be both costly and time-consuming, particularly when training agents for complex driving tasks. This not only limits scalability but also makes real-world deployment more challenging [10]. Second, DRL models generally lack transparency and interpretability, which impedes their ability to make reliable decisions in rare or out-of-distribution scenarios [20]. This lack of transparency makes DRL less reliable for safety-critical applications, such as handling unexpected or unfamiliar traffic situations.
To address these issues, the integration of expert knowledge through Reinforcement Learning from Human Feedback (RLHF) has been proposed [21]. RLHF allows for faster convergence and improved robustness by leveraging human expertise to guide the learning process, reducing the number of required interactions with the environment. However, RLHF comes with its own set of challenges. First, it is resource-intensive due to the need for extensive human annotations [22]. Additionally, the human feedback may not cover the full range of possible driving scenarios, limiting the agent’s ability to generalize effectively to unseen situations. These limitations point to the need for a more efficient and scalable method that integrates expert guidance while addressing DRL’s inherent drawbacks.
@@ II-B LLMs in Decision-Making
LLMs have shown considerable promise in various high-level decision-making tasks, including autonomous driving. LLMs, such as GPT-4, have demonstrated their ability to handle complex reasoning, interpretation, and contextual awareness [23]. For example, LanguageMPC [24] leverages the common-sense reasoning capabilities of LLMs to guide Model Predictive Control (MPC) parameters for autonomous vehicles. Similarly, Fu et al. [25] and Wen et al. [26] have explored the application of LLM-based reasoning, interpretation, and memory capabilities to assist autonomous decision-making, particularly in complex and dynamic traffic environments. These models help in interpreting driving scenarios and proposing context-aware strategies based on learned knowledge.
Despite these promising developments, the practical deployment of LLMs in autonomous driving faces several limitations. One of the key challenges is the high computational cost associated with running LLMs in real-time, making it difficult to meet the responsiveness required for safety-critical applications [13]. Additionally, LLMs typically generate outputs with a degree of randomness, which can result in unpredictable actions that are unsuitable for tasks demanding consistent and reliable decision-making. This unpredictability is particularly problematic in autonomous driving, where even minor deviations from expected behavior can have serious safety implications. Thus, while LLMs offer significant potential for decision-making in autonomous driving, their practical use as standalone decision-making agents is limited by their real-time performance and output consistency.
II-C Hybrid DRL-LLM Approaches
With the rapid development of LLMs and DRL in various fields, researchers are increasingly exploring the synergistic potential of combining these two paradigms. While numerous studies have focused on using DRL methods to optimize and fine-tune LLMs to enhance their generative capabilities and task adaptability, the utilization of LLMs to assist DRL remains relatively underexplored, particularly in the context of autonomous driving decision-making.
Existing research has begun to investigate how the reasoning and knowledge capabilities of LLMs can improve the exploration efficiency and learning effectiveness of RL agents[27]. For example, Zhang et al.[28] developed a semi-parametric RL framework based on LLMs by configuring long-term memory modules; Similarly, Trirat et al.[29] employed LLMs to achieve full-process automated machine learning, while Ma et al.[30] realized the automatic design of reward functions in RL without requiring specific enhancement tasks. Despite these advancements, the environmental understanding capabilities of LLMs are still not fully leveraged, and effective integration between LLMs and RL remains a challenge. Current approaches lack a comprehensive methodology for combining the strengths of both LLMs and RL, resulting in an under-utilized potential to improve decision-making processes in autonomous driving systems.
@@ III-2 Action space
This work focuses on leveraging LLMs to provide high-level guidance for DRL, rather than controlling low-level vehicle dynamics. Consequently, the action space $\mathcal{A}$ consists of five high-level maneuvers:
$\mathcal{A}=\{slowdown,cruise,speedup,turnleft,turnright\}$ (3)
Once a high-level maneuver is chosen, the corresponding steering and throttle commands are generated by a lower-level PID controller, enabling the vehicle to execute lateral and longitudinal movements.
@@ IV-A Framework overview
TeLL-Drive leverages the prior knowledge of LLMs to guide the exploration and learning of DRL agents in diverse, complex traffic scenarios. By introducing policy integration, TeLL-Drive enhances sample efficiency and optimizes learning outcomes. As illustrated in Fig. 2, the framework comprises two main components: the LLM Teacher and the DRL Student. Based on multi-module collaboration, the Teacher Agent generates robust decision through its three key modules: Decision Engine, which provides real-time guidance; Memory Repository, which stores past experiences for context; and Reflective Evaluator, which refines the guidance based on previous performance. While the Student Agent refines the Teacher’s actions through a multi-head attention-based policy-integration mechanism, integrating its own exploration experiences to effectively acquire knowledge from the LLM and enhance learning efficiency and quality.
@@ IV-B1 Decision Engine
Building on these real-time and historical insights, the engine constructs a comprehensive prompt $\mathcal{P}_{t}$ that integrates TTCP-derived risks, scenario-specific experiences, and traffic knowledge. This prompt includes road geometry, vehicle positions, and conflict zones, along with domain heuristics to provide the LLM with a rich contextual foundation for action proposals. To enhance reliability and minimize hallucinations, we then employ a chain-of-thought approach  [31] in which the model iteratively evaluates collision severity, short- and long-term maneuver consequences, and broader traffic implications. This structured reasoning process reduces logical inconsistencies, resulting in safer and more interpretable autonomous driving policies.
@@ IV-B2 Memory Repository
$\mathcal{P}_{t}\leftarrow\text{Retrieval}\Bigl{(}\mathbf{s}_{t},\mathcal{M}\Bigr{)}$ (6)
@@ IV-B3 Reflective Evaluator
where $\tau_{TTCP}(s_{t},a_{t})$ is the TTCP for action $a_{t}$ in state $s_{t}$, $\mathbb{I}\{\cdot\}$ is an indicator function for specific infractions, and $\beta$ is a weighting constant.Any episode with $\max_{t}\Omega(s_{t},a_{t})\geq\delta$ is flagged for further reflection.
@@ IV-C1 Actor-Critic Algorithm with Policy Constraint
An actor-critic framework[32] is adopted, where both the state-value function $V^{\pi}$ and the action-value function $Q^{\pi}$ are recursively estimated. For a policy $\pi(\mathbf{a}\mid\mathbf{s})$, the state-value function and the corresponding action-value function at state $\mathbf{s}_{t}$ is:
@@ IV-C2 Policy Distillation and Fusion
$\hat{\pi}^{T}(\mathbf{a}\mid\mathbf{s}_{t})$ $=softmax\bigl{(}\mathcal{W}_{p}\,\mathbf{h}_{t}^{T}+\mathbf{b}_{p}\bigr{)}$ (15)
$\hat{Q}^{T}(\mathbf{s}_{t},\mathbf{a})$ $=\mathcal{W}_{q}\,\mathbf{h}_{t}^{T}+\mathbf{b}_{q}$ (16)
@@
$\mathbf{h}_{t}=\alpha_{t}\,V+\mathbf{h}_{t}^{S}=\alpha_{t}\,\mathcal{W}_{V}\,\mathbf{h}_{t}^{T}+\mathbf{h}_{t}^{S}$ (19)
@@
$\tilde{\pi}(\mathbf{a}\mid\mathbf{s}_{t})$ $=softmax\bigl{(}\mathbf{W}_{\pi}\,\mathbf{h}_{t}+\mathbf{b}_{\pi}\bigr{)}$ (20)
$\tilde{Q}(\mathbf{s}_{t},\mathbf{a})$ $=\mathbf{W}_{Q}\,\mathbf{h}_{t}+\mathbf{b}_{Q}$ (21)
@@
$\mathcal{L}_{\mathrm{distill}}=-\,\mathbb{E}_{(s,\mathbf{a}_{T})\in\mathcal{D}_{T}}\bigl{[}\log\,\hat{\pi}^{T}(\mathbf{a}_{T}\mid s)\bigr{]}$ (22)
@@
V Simulation and Performance Evaluation
@@ V-B Implementation Details
Both our model and the baseline methods utilize a policy network composed of a multilayer perceptron (MLP) with two hidden layers of size $128\times 128$. We employ two self-attention heads, each also of dimension 128, to fuse the Student and Teacher representations.The clip range is dynamically adjusted using a linear schedule, starting with an initial value and decreasing according to the remaining training progress. Each model is trained for at least $10^{5}$ time steps, with an evaluation performed every 500 time steps. We use GPT-4o-mini[11] as our LLM backbone, which shows reliable logical reasoning and real-time decision-making in driving tasks; it serves as the Teacher Agent for only the first 10% of training steps, after which constraints are gradually relaxed to encourage independent exploration. The specific parameter settings are shown in Table I. All experiments run on a computing platform equipped with Intel(R) Core(TM) i7-14700K CPU, an NVIDIA GeForce RTX 4080 SUPER GPU and 32 GB of RAM.
TABLE I: Hyperparameters used in the experiment
Symbol Meaning Value
$\alpha$ Learning rate $5\times 10^{-4}$
$\mathcal{N}_{train}$ Minimum total training steps $1\times 10^{5}$
$\gamma$ Discount factor 0.99
$\epsilon$ Initial value of clip range 0.2
$\mathcal{B}$ Training batch size 128
$\mathcal{N}_{B}$ Rollout buffer size 1600
$\|\mathcal{M}\|$ Capacity of Memory Repository 20
$\mathcal{N}_{shot}$ Number of examples for few-shot learning 3